OpenStudy (precal):

I have to find the horizontal tangent and vertical tangent for the following function

OpenStudy (precal):

\[x^3+y^3=4xy+1\]

OpenStudy (precal):

for the horizontal tangent, I know I need to take the derivative and solve for 0

OpenStudy (precal):

Not sure the approach for the vertical tangent

OpenStudy (anonymous):

By horizontal and vertical, do we mean tangent and normal?

OpenStudy (precal):

For what values of x will the curve x^3+y^3=4xy+1 have a horizontal tangent?

OpenStudy (precal):

ok I thought I had the right approach, not sure anymore

hartnn (hartnn):

horizontal tangent : dy/dx = 0 vertical tangent : dx/dy = 0

OpenStudy (precal):

\[\frac{ dy }{ dx }=\frac{ 4xy-3x^2 }{ 3y^2-1 }\]

OpenStudy (precal):

how do I solve this one for 0?

OpenStudy (precal):

\[4xy-3x^2=0\]

OpenStudy (tkhunny):

Careful, if the denominator is also zero, you may not have what you think.

OpenStudy (ikram002p):

by horizontal tangent switch the equation

OpenStudy (precal):

do you mean for the vertical tangent switch the equations? \[\frac{ dx }{ dy }=\frac{ 3y^2-1 }{ 4xy-3x^2 }\] I can solve this one, \[3y^2-1=0\] \[y=\pm \sqrt{\frac{ 1 }{ 3 }}\]

hartnn (hartnn):

so you got the y co-ordinate. get the x co-ordinate from the original equation. because vertical lines have x value constant (hence their equation is x=c)

OpenStudy (precal):

so do I sub this y value into the original to find the x values

hartnn (hartnn):

yes

OpenStudy (zarkon):

you might want to double check your derivative

OpenStudy (precal):

ok still don't know what to do with my horizontal tangent ok I will double check it

OpenStudy (precal):

oh I dropped something, I will redo it

hartnn (hartnn):

you're not given any point, at which the tangents are to be found ?

OpenStudy (precal):

\[\frac{ dy }{ dx }=\frac{ 4xy-3x^2 }{ 3y^2-4x}\]

OpenStudy (precal):

ok this is the correct derivative

OpenStudy (precal):

No, this is what I was given..... For what values of x will the curve \[x^3+y^3=4xy+1\] have a horizontal tangent? Show your work and explain your thinking.

OpenStudy (precal):

My second part, In terms of y, describe the values of x for which the curve \[x^3+y^3=4xy+1\] will have a vertical tangent? Show your work and explain your thinking.

OpenStudy (precal):

so for this part (the second part) if I solve for y I get: \[y=\pm \sqrt{\frac{ 4x }{ 3 }}\]

hartnn (hartnn):

what values of x in terms of y so, x= 3y^2/4

OpenStudy (zarkon):

are you sure that derivative is correct

OpenStudy (precal):

omg am I that dumb???? ok I will triple check it.......my brain is fried

OpenStudy (precal):

ok here is my last attempt......\[\frac{ dy }{ dx }=\frac{ 4y-3x^2 }{ 3y^2-4x }\]

hartnn (hartnn):

you're learning, silly mistakes will happen

OpenStudy (zarkon):

that is correct

OpenStudy (precal):

ok did I get it correct this time? No, I am just tired (been doing math all day long) Maybe I have reached my limit

OpenStudy (precal):

haha calculus joke

OpenStudy (precal):

ok my second part is ok is that what they meant for the second part

OpenStudy (precal):

1st part am I solving for the solution in terms of x?

OpenStudy (precal):

so is the first part \[x=\pm \sqrt{\frac{ 4y }{ 3 }}\]

OpenStudy (precal):

not sure anymore......

hartnn (hartnn):

plug that into original equation

OpenStudy (precal):

Is that what I need to do for both of them?

hartnn (hartnn):

you will get equation only in 'x' solving that equation will give you x values for which you have horizontal tangent

hartnn (hartnn):

for 2nd part, you CAN do the same thing, but it is not asked for "In terms of y" x =3y^2/4 is good enough

OpenStudy (precal):

ok thanks, I am not sure this is what I need to do....I don't recall these equations being sooooooo bad or so complex.....